# Operating System And Artificial Intelligence: A Systematic Review

### v1

@@ 10 Conclusion
The synthesis of AI and OS is reshaping the landscape of contemporary computing, as evidenced by our meticulous review of 108 primary studies. AI’s integration into OS frameworks has catalyzed significant improvements in key areas like memory oversight, task coordination, and security vigilance. By leveraging intelligent decision-making beyond traditional rule-bound paradigms, AI bolsters resource efficiency and system agility.
In tandem, advanced OS designs are morphing to better serve AI applications, especially as data-intensive workloads surge across various computing platforms. The OS’s evolving role in adeptly handling specialized hardware and optimizing data flow underscores its importance in sustaining AI software performance.
The integration of LLMs into OS interfaces, as highlighted in our analysis of 17 studies, marks a pivotal advancement in user experience. Conversational interfaces enable seamless human-computer interaction, simplifying device management and personalizing user assistance through natural language processing.
Despite the optimistic outlook, integrating AI into OS presents critical challenges. Safeguarding against security breaches and privacy infringements while maintaining efficient AI operations necessitates rigorous design considerations. Moreover, the demand for streamlined AI algorithms that can operate within constrained environments remains a high-priority research area.
Future research agendas should concentrate on the domain of intelligent operating systems, capitalizing on the transformative potential they offer. Central to this endeavor is the exploration of LibOS. With its modular structure, LibOS facilitates the seamless incorporation of AI enhancements, supports granular control, fosters innovation, and upholds stringent security standards. LibOS stands as a robust foundation for the evolution of intelligent computing. Thus, advancing LibOS technologies should be a focal point for researchers aiming to realize the full scope of intelligent OS capabilities and drive the computing industry toward greater adaptability, efficiency, and resilience.
Acknowledgments
@@
This work was supported in part by the Major Program of National Natural Science Foundation of Zhejiang(LD24F020014), and in part by the Zhejiang Pioneer (Jianbing) Project (2024C01032), and in part by the Key R&D Program of Ningbo(2023Z235), and in part by the Ningbo Yongjiang Talent Programme(2023A-198-G).

### v2

@@ 9 Conclusion
In summary, this review has comprehensively explored the current state of AI-OS integration, highlighting significant advancements and identifying areas for further investigation. Key findings include the effective utilization of AI techniques, such as ML and large language models, to improve OS functionalities like memory management, process scheduling, and intrusion detection. The review underscores the importance of secure and efficient AI integration within OS frameworks, emphasizing the need for robust validation methods and evaluation metrics. Additionally, the potential of novel OS architectures that leverage AI capabilities is outlined, paving the way for intelligent, modular, adaptable, and self-aware operating systems. Future research should focus on addressing the challenges associated with integrating AI into OSs, including the development of more efficient training strategies, the creation of larger and more diverse datasets, and the improvement of AI models’ interpretability and explainability. Exploring the ethical implications of AI in OSs is also crucial, ensuring that the benefits of these technologies are balanced against potential risks. The convergence of AI and OS represents a pivotal arena for innovation in computing technology, with the development of intelligent OSs critical in shaping the future of computing.
Acknowledgment
@@
This work was supported in part by the National Science Foundation of China under Grants (62472375,62125206), and in part by the Major Program of National Natural Science Foundation of Zhejiang (LD24F020014), and in part by the Zhejiang Pioneer (Jianbing) Project (2024C01032), and in part by the Key R&D Program of Ningbo (2023Z235), and in part by the Ningbo Yongjiang Talent Programme (2023A-198-G), and in part by the fund of Laboratory for Advanced Computing and Intelligence Engineering.
A Tables
Resource name Search string in Initial selection
ACM Digital Library Abstract 210
IEEE Xplore Digital Library Abstract 279
arXiv by Cornell University Abstract 158
Google Scholar Title 177
Table 1: Search results from different data sources.
Abbreviation Description
MLP Multi-Layer Perceptron
RNN Recurrent Neural Network
CNN Convolutional Neural Network
LSTM Long Short-Term Memory
DT Decision Tree
RF Random Forest
GA Genetic algorithm
LLM Large Language Model
DNN Deep Neural Networks
SDN Software Defined Network
LRU least recently used
Table 2: Abbreviations
B Detailed information on selected papers
Figure 11 depicts the distribution of publication years for research. The figure indicates that from 2019 to 2021, the number of published works was relatively low and stable. However, starting from the year 2023, there has been a steady increase in the number of publications. This could reflect an increasing level of activity in this field of research, as well as potential trends and growing interest in the subject.
Figure 12(a) illustrates the distribution of publication venues for research. Notably, the ACM International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS) has accepted a significantly greater number of publications.
The distribution of countries is shown in Figure 12(b), where the USA, China, and India are the top three contributors to the total number of publications.
Figure 11: Year Distribution of Paper Published
(a) Venue Distribution of Paper Published
(b) Country Distribution of Paper Published
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used Qianwen and ChatGPT in order to improve the readability in the writing process. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.